# Conformalized Bayesian Inference, with Applications to Random Partition Models

### v1

@@ Abstract
Bayesian posterior distributions naturally represent parameter uncertainty informed by data. However, when the parameter space is complex, as in many nonparametric settings where it is infinite dimensional or combinatorially large, standard summaries such as posterior means, credible intervals, or simple notions of multimodality are often unavailable, hindering interpretable posterior uncertainty quantification. We introduce Conformalized Bayesian Inference (CBI), a broadly applicable and computationally efficient framework for posterior inference on nonstandard parameter spaces. CBI yields a point estimate, a credible region with assumption-free posterior coverage guarantees, and a principled analysis of posterior multimodality, requiring only Monte Carlo samples from the posterior and a notion of discrepancy between parameters. The method builds a discrepancy-based kernel density score for each parameter value, yielding a maximum-a-posteriori-like point estimate and a credible region derived from conformal prediction principles. The key conceptual step underlying this construction is the reinterpretation of posterior inference as prediction on the parameter space. A final density-based clustering step identifies representative posterior modes. We investigate a number of theoretical and methodological properties of CBI and demonstrate its practicality, scalability, and versatility in simulated and real data clustering applications with Bayesian random partition models. [^1] The authors are grateful to Michele Guindani, Peter Müller and Stephen G. Walker for their insightful comments, and to Yunshan Duan, Ziyi Song and Wenyi Wang for sharing their model outputs used for the real data applications presented in the paper. Nicola Bariletto gratefully acknowledges funding from the “G. Mortara” scholarship by the Bank of Italy.
@@ 1 Introduction
Bayesian statistical methods have become a standard tool in probabilistic modeling, primarily due to their natural ability to represent uncertainty in the phenomena under study. In the Bayesian framework, the generative process underlying the observed data is formalized as a statistical model with parameters to be inferred from the data. Uncertainty about these parameters is expressed through a prior distribution, which is updated to the posterior distribution given the data, and subsequently to the predictive distribution for unseen observations. A central advantage of the Bayesian framework is that, when inference is required over a parameter space $\Theta$ indexing the statistical model, it is enough to specify a prior probability distribution on $\Theta$ and to compute, often approximately (for instance, via Monte Carlo methods), the corresponding posterior distribution. This provides a coherent basis for point estimation—selecting a single point that summarizes the posterior distribution—and uncertainty quantification—characterizing the regions of $\Theta$ on which the posterior places significant mass. This paradigm has proven particularly powerful in the nonparametric setting, where modeling flexibility is achieved by allowing $\Theta$ to be a complex, and often infinite-dimensional, parameter space.
@@
This abstract flexibility, however, often comes at the price of some practical difficulties. While parameter spaces associated with flexible priors can capture complex aspects of the data-generating process, they may lack a conventional structure (for example, a full vector space or ordering structure) that allows for an interpretable summary of the posterior distribution through quantities such as a posterior mean/mode or credible interval/band. This challenge is pervasive. Consider, for instance, models whose parameter spaces consist of data partitions (Wade,, 2023), covariance matrices (Leonard and Hsu,, 1992; Yang and Berger,, 1994), graphs (Nowicki and Snijders,, 2001; Orbanz and Roy,, 2014), mixing measures (Nguyen,, 2013), among innumerable other examples. In such cases, although Monte Carlo methods, often in the form of Markov Chain Monte Carlo (shortened to MCMC, Metropolis et al.,, 1953; Hastings,, 1970; Tierney,, 1994; Gamerman and Lopes,, 2006), provide good empirical approximations of the posterior distribution, the complexity of the sampled objects hinders interpretable inference. This raises a few fundamental questions: What constitutes a suitable point estimate of the random parameter $\theta\in\Theta$ under the posterior distribution? How can one define a credible region with reasonable size and a prescribed posterior coverage level? Is the posterior distribution concentrated in a single, homogeneous region of $\Theta$, or does it assign substantial mass to multiple distinct regions? When only Monte Carlo samples are available and the parameter space lacks a familiar Euclidean-like structure, no general or obvious answers exist to these questions.
@@
For specific instances of complex parameter spaces, these challenges have motivated a variety of solutions, ranging from theoretically grounded to ad-hoc. For example, in point estimation, a common decision-theoretic approach specifies a loss function on $\Theta\times\Theta$ and selects the minimizer of the posterior expected loss, typically approximated through Monte Carlo sampling (Bernardo and Smith,, 1994). In contrast, questions concerning credible regions and posterior multimodality are often addressed through problem-specific heuristics that exploit the structure of $\Theta$. For instance, when $\Theta$ denotes the space of data clusterings arising from random partition models, several tailored approaches have been proposed (Wade and Ghahramani,, 2018; Balocchi and Wade,, 2025); since random partition models will serve as our main illustrative application, we defer a detailed discussion of these methods. A useful strategy is to examine low-dimensional aspects of the parameter (Woody et al.,, 2021; Bolfarine et al.,, 2025), but while this can offer preliminary insight into posterior uncertainty, it risks discarding the very flexibility afforded by using a rich parameter space $\Theta$. It is therefore desirable to develop general and principled frameworks for addressing these questions simultaneously, coherently, and efficiently, given only Monte Carlo samples from a posterior distribution over $\Theta$.
@@
In this article, we make a step in this direction by introducing a broadly applicable methodology grounded in the theories of conformal prediction (Vovk et al.,, 2005; Shafer and Vovk,, 2008; Angelopoulos and Bates,, 2023) and density-based clustering (Rodriguez and Laio,, 2014; Ester et al.,, 1996; Kriegel et al.,, 2011), which we term Conformalized Bayesian Inference (CBI). The proposed pipeline is as follows. Assume that, given a posterior distribution $\Pi$ on $\Theta$, we have at our disposal $T\in\mathbb{N}$ independent and identically distributed (iid) Monte Carlo samples $\bm{\theta}=\{\theta_{1},\dots,\theta_{T}\}$, each marginally distributed according to $\Pi$. The iid assumption may be either true exactly by the nature of the adopted Monte Carlo scheme, or can be though to hold approximately in the case of MCMC samples, as long as the sampler chain has been properly thinned (see Theorem 2 below). We then split the samples into two subsets: a training set $\bm{\theta}^{1}=\{\theta_{1},\dots,\theta_{S}\}$ and a calibration set $\bm{\theta}^{2}=\{\theta_{S+1},\dots,\theta_{T}\}$. The key idea is to compute, for each calibration sample $\theta\in\bm{\theta}^{2}$, a score
$\theta\mapsto s(\theta;\bm{\theta}^{1})\in\mathbb{R},$
which depends on the training set and is designed to capture a notion of likelihood or density of $\theta$ under the posterior distribution. Once the scores are available, the CBI framework proceeds through the following three main steps:
@@
    $\theta_{\star}\in\arg\max_{\theta\in\bm{\theta}^{2}}s(\theta;\bm{\theta}^{1}).$
    This may be interpreted as a maximum-a-posteriori-like estimator, provided that the score function has been constructed to reflect some notion of likelihood under $\Pi$.
  • Step 2: Credible region. By exchangeability of the calibration scores, standard conformal prediction arguments yield a credible region $\mathcal{C}_{1-\alpha}(\bm{\theta}^{2})\subseteq\Theta$, based on the quantiles of $\{s(\theta;\bm{\theta}^{1}):\theta\in\bm{\theta}^{2}\}$, having a prescribed $(1-\alpha)\times 100\%$ coverage under $\Pi$ (for $\alpha\in(0,1)$, in a precise statistical sense to be defined later). The key conceptual step leading to this construction is to interpret inference based on posterior Monte Carlo samples as prediction on the parameter space according to the posterior distribution, and to leverage conformal prediction tools accordingly. Two appealing features of this procedure are that (i) the attached coverage guarantee is independent of any feature of the target posterior $\Pi$, accommodating parameter spaces of arbitrary dimension and with almost no structure, and (ii) for a new sample $\theta$, it provides a simple way to define a standardized measure of typicality $\hat{p}(\theta;\bm{\theta}^{2})\in[0,1]$ under $\Pi$, which may also be formally interpreted as a $p$-value under an intuitive distributional null hypothesis.
  • Step 3: Multimodality analysis. Finally, the interpretation of $s(\cdot;\bm{\theta}^{1})$ as a proxy for posterior density allows one to explore the possibly multimodal structure of the posterior distribution through density-based clustering techniques. By equipping $\Theta$ with a meaningful discrepancy or metric $\mathcal{D}$, one can identify distinct high-density regions, thereby revealing whether the posterior distribution places substantial mass around multiple, well-separated parameter values.
@@
While CBI can be described in full generality without reference to any particular statistical model, it is useful to ground the discussion in a concrete setting involving a parameter space $\Theta$ of interest. Therefore, before presenting a general treatment of CBI, in Section 2 we introduce a fundamental class of models, namely random partition models, on which the strengths of our methodology can be clearly demonstrated. An additional advantage of working with random partition models is the existence of a rich literature addressing point estimation and uncertainty quantification based on posterior Monte Carlo samples over the space of data partitions (Wade and Ghahramani,, 2018; Wade,, 2023; Balocchi and Wade,, 2025), providing a natural conceptual benchmark for our discussion. Moreover, since clustering is an inherently difficult problem (Hennig,, 2015), Bayesian posteriors over data partitions often display substantial uncertainty, frequently in the form of multimodality, which our framework is specifically designed to handle. In Section 3 we then describe CBI in detail and illustrate it with simulated data analyzed with random partition models. [^2] However, see the Supplementary Material for two additional experiments applying CBI to models where the parameter space consists of mixing measures endowed with the Wasserstein-1 distance (Nguyen,, 2013; Villani et al.,, 2008) and covariance matrices endowed with the operator norm distance. Section 4 showcases three real-world applications of CBI to Bayesian clustering of (i) the classical Galaxy velocities dataset (Roeder,, 1990), (ii) colorectal cancer spatial transcriptomics data (Lee et al.,, 2020; Duan et al.,, 2025), and (iii) event-related potential (ERP) waveform functional data (Song et al.,, 2025; Kappenman et al.,, 2021). Section 5 concludes the article with a discussion of the current limitations and future directions related to CBI.
@@ 2 Motivation: random partition models
allows for coherent inference on the clustering structure of the observed data. In practice, inference on clustering may be formulated at the latent level of a hierarchical density mixture model, where heterogeneity among subpopulations is interpreted as observations being generated from different components of a mixture model (Fruhwirth-Schnatter et al.,, 2019; Lo,, 1984; Escobar and West,, 1995; Lijoi et al.,, 2005; Barrios et al.,, 2013; Lijoi et al.,, 2020), or, more recently, through approaches based on tailored losses (Rigon et al.,, 2023) or level sets (Buch et al.,, 2024). We refer the reader to Wade, (2023), Grazian, (2023), and the review section of Balocchi and Wade, (2025) for comprehensive and up-to-date overviews of the field.
@@
The combinatorially large size of $\Theta$ typically prevents exact posterior evaluation, and approximate computational methods, most often in the form of MCMC sampling, are adopted. For instance, if the random partition model is specified within a hierarchical mixture framework, standard samplers (Escobar and West,, 1995, 1998; MacEachern and Müller,, 1998; Favaro and Teh,, 2013; Neal,, 2000) will, at each iteration $t=1,\ldots,T$, draw a sample $\theta_{t}\sim\Pi$ (in addition to sampling the other mixture parameters). These samples are usually stored as vectors of cluster assignments $C_{t}=(c_{t1},\ldots,c_{tn})$, encoding the partition (equivalence) relation $\sim_{\theta_{t}}$ via $X_{i}\sim_{\theta_{t}}X_{j}$ if and only if $c_{ti}=c_{tj}$, for all $i,j=1,\ldots,n$. Assuming appropriate thinning of the chain and a warm start $\theta_{1}\sim\Pi$, the resulting samples $\theta_{1},\ldots,\theta_{T}$ will serve as the basis of our CBI procedure.
@@
A widely used metric on the space of partitions $\Theta$ is the Variation-of-Information (VI) metric (Meilă,, 2007), denoted by $\mathcal{D}_{VI}$ and defined as follows. Take any two partitions $\theta,\theta^{\prime}\in\Theta$ with associated cluster assignment vectors $C,C^{\prime}$ and respective numbers of clusters $K$ and $K^{\prime}$. Let $\{\bar{c}_{1},\ldots,\bar{c}_{K}\}$ and $\{\bar{c}^{\prime}_{1},\ldots,\bar{c}^{\prime}_{K^{\prime}}\}$ denote the sets of unique cluster labels in each vector, and define $n_{j,*}:=|\{i=1,\ldots,n:c_{i}=\bar{c}_{j}\}|$, $n_{*,k}:=|\{i=1,\ldots,n:c^{\prime}_{i}=\bar{c}^{\prime}_{k}\}|$, and $n_{j,k}:=|\{i=1,\ldots,n:c_{i}=\bar{c}_{j},\,c^{\prime}_{i}=\bar{c}^{\prime}_{k}\}|$, for all $j=1,\ldots,K$ and $k=1,\ldots,K^{\prime}$. Then,
$\mathcal{D}_{VI}(\theta,\theta^{\prime})$ $:=$
$-\,\sum_{j=1}^{K}$ $\frac{n_{j,*}}{n}\log_{2}\left(\frac{n_{j,*}}{n}\right)\,-\,\sum_{k=1}^{K^{\prime}}\frac{n_{*,k}}{n}\log_{2}\left(\frac{n_{*,k}}{n}\right)\,-\,2\sum_{j=1}^{K}\sum_{k=1}^{K^{\prime}}\frac{n_{j,k}}{n}\log_{2}\left(\frac{n\cdot n_{j,k}}{n_{j,*}\cdot n_{*,k}}\right).$ (1)
@@
Meilă, (2007) showed that $\mathcal{D}_{VI}$ is a proper metric on $\Theta$, and that it enjoys several appealing theoretical properties, such as its information-theoretic interpretation in terms of self-entropy and cross-entropy between partitions, while also being efficiently computable. Although alternative discrepancy measures have been proposed, such as the Binder loss (Binder,, 1978; Dahl,, 2006; Lau and Green,, 2007), the adjusted Rand index (Fritsch and Ickstadt,, 2009), the normalized VI (Vinh et al.,, 2010; Rastelli and Friel,, 2018), and more (Quintana and Iglesias,, 2003; Dahl et al.,, 2022; Nguyen and Mueller,, 2024), we focus on the VI metric in what follows, given its widespread use and well-established role as the standard metric in Bayesian clustering.
@@
The large size of $\Theta$ and its discrete nature make it an ideal example of a setting where point estimation and posterior uncertainty quantification are highly non-trivial, even when Monte Carlo samples from $\Pi$ are available. The popularity of clustering and the success of Bayesian approaches in this domain have given rise to an extensive literature on the topic. A seminal contribution is due to Wade and Ghahramani, (2018), who popularized the use of the VI metric in Bayesian clustering and proposed to address point estimation via posterior expected VI minimization, while constructing (approximations of) VI credible balls with prescribed posterior mass for uncertainty quantification. Computational and theoretical extensions of this framework, often but not exclusively focused on point estimation, include Rastelli and Friel, (2018); Dahl et al., (2022); Nguyen and Mueller, (2024); Buch et al., (2024), among others. A recent and insightful perspective was introduced by Balocchi and Wade, (2025), who emphasized the importance of accounting for multimodality in the posterior distribution over $\Theta$ and proposed summarizing $\Pi$ with the closest $L$-atom mixture of point masses in the Wasserstein-over-VI metric, yielding a fixed number $L$ of representative partitions corresponding to the distinct posterior modes.
@@
3 Methodology: Conformalized Bayesian Inference
Recall the problem we aim to solve: given iid Monte Carlo samples $\bm{\theta}$ drawn from the parameter space $\Theta$ and split into a training subset $\bm{\theta}^{1}$ and a calibration subset $\bm{\theta}^{2}$, how can we perform point estimation and uncertainty quantification in practice? Following the pipeline described in the introduction, the first key step is to define a scoring function $\theta\mapsto s(\theta;\bm{\theta}^{1})$ that captures a notion of density under the posterior distribution. We propose a simple and robust solution.

### v2

@@ Abstract
Bayesian posterior distributions naturally represent parameter uncertainty informed by data. However, when the parameter space is complex, as in many nonparametric settings where it is infinite-dimensional or combinatorially large, standard summaries such as posterior means, credible intervals, or simple notions of multimodality are often unavailable, hindering interpretable posterior uncertainty quantification. We introduce Conformalized Bayesian Inference (CBI), a broadly applicable and computationally efficient framework for posterior inference on nonstandard parameter spaces. CBI yields a point estimate, a credible region with assumption-free posterior coverage guarantees, and a principled analysis of posterior multimodality, requiring only Monte Carlo samples from the posterior and a notion of discrepancy between parameters. The method builds a pseudo-density score for each parameter value, yielding a MAP-like point estimate and a credible region derived from conformal prediction principles. The key conceptual step underlying this construction is the reinterpretation of posterior inference as prediction on the parameter space. A final density-based clustering step identifies representative posterior modes. We investigate a number of theoretical and methodological properties of CBI and demonstrate its practicality, scalability, and versatility in simulated and real data clustering applications with random partition models. An accompanying Python library, cbi_partitions, is available at https://github.com/nbariletto/cbi_partitions_repo. [^1] A step-by-step tutorial is found at https://nbariletto.github.io/cbi_partitions_repo/.
Keywords: Bayesian Clustering, Bayesian Nonparametrics, Conformal Prediction, Monte Carlo Methods, Multimodality.
@@ 1 Introduction
Bayesian statistical methods have become a standard tool in probabilistic modeling, primarily due to their natural ability to represent uncertainty in the phenomena under study. In the Bayesian framework, the generative process underlying the observed data is formalized as a statistical model with parameters to be inferred from the data. Uncertainty about these parameters is expressed through a prior distribution, which is updated to the posterior distribution given the data, and subsequently to the predictive distribution for unseen observations. A central advantage of the Bayesian framework is that, when inference is required over a parameter space $\Theta$ indexing the statistical model, it is enough to specify a prior probability distribution on $\Theta$ and to compute, often approximately (for instance, via Monte Carlo methods), the corresponding posterior distribution. This provides a coherent basis for point estimation—selecting a single parameter value that summarizes the posterior distribution—and uncertainty quantification—characterizing the regions of $\Theta$ on which the posterior places significant mass. This paradigm has proven particularly powerful in the nonparametric setting, where modeling flexibility is achieved by allowing $\Theta$ to be a complex, and often infinite-dimensional, parameter space.
@@
This abstract flexibility, however, often comes at the price of some practical difficulties. While parameter spaces associated with flexible priors can capture complex aspects of the data-generating process, they may lack a conventional structure (for example, a vector or ordering structure) that allows for an interpretable summary of the posterior distribution through quantities such as a posterior mean/mode or a credible interval/band. This challenge is pervasive. Consider, for instance, models whose parameter spaces consist of data partitions (wade2023bayesian), covariance matrices (leonard1992bayesian, yang1994estimation), graphs (nowicki2001estimation, orbanz2014bayesian), mixing measures (nguyen2013convergence), among innumerable other examples. In such cases, although Monte Carlo methods, often in the form of Markov Chain Monte Carlo (shortened to MCMC, metropolis1953equation, hastings1970monte, tierney1994markov, gamerman2006markov), provide good empirical approximations of the posterior distribution, the complexity of the sampled objects hinders interpretable inference. This raises a few fundamental questions: What constitutes a suitable point estimate of the random parameter $\theta\in\Theta$ under the posterior distribution? How can one define a credible region with reasonable size and a prescribed posterior coverage level? Is the posterior distribution concentrated in a single, homogeneous region of $\Theta$, or does it assign substantial mass to multiple distinct regions? When only Monte Carlo samples are available and the parameter space lacks a familiar Euclidean-like structure, no general or obvious answers exist to these questions.
@@
For specific instances of complex parameter spaces, these challenges have motivated a variety of solutions, ranging from theoretically grounded to ad hoc. For example, in point estimation, a common decision-theoretic approach specifies a loss function on $\Theta\times\Theta$ and selects the minimizer of the posterior expected loss, typically approximated through Monte Carlo sampling (bernardo1994bayesian). In contrast, questions concerning credible regions and posterior multimodality are often addressed through problem-specific heuristics that exploit the structure of $\Theta$. For instance, when $\Theta$ denotes the space of data clusterings arising from random partition models, several tailored approaches have been proposed (wade2018clustering, balocchi2025understanding); since random partition models will serve as our main illustrative application, we defer a detailed discussion of these methods. A useful strategy is to examine low-dimensional aspects of the parameter (woody2021model, bolfarine2025lower), but while this can offer preliminary insight into posterior uncertainty, it risks discarding the very flexibility afforded by using a rich parameter space $\Theta$. It is therefore desirable to develop general and principled frameworks for addressing these questions simultaneously, coherently, and efficiently, given only Monte Carlo samples from a posterior distribution over $\Theta$.
@@
In this article, we make a step in this direction by introducing a broadly applicable methodology grounded in the theories of conformal prediction (vovk2005algorithmic, shafer2008tutorial, angelopoulos2023conformal) and density-based clustering (rodriguez2014clustering, ester1996density, kriegel2011density), which we term Conformalized Bayesian Inference (CBI). The proposed pipeline is as follows. Assume that, given a posterior distribution $\Pi$ on $\Theta$, we have at our disposal $T\in\mathbb{N}$ independent and identically distributed (iid) Monte Carlo samples $\boldsymbol{\theta}=\{\theta_{1},\dots,\theta_{T}\}$, each marginally distributed according to $\Pi$. The iid assumption may be either true exactly by the nature of the adopted Monte Carlo scheme, or can be thought to hold approximately in the case of MCMC samples, as long as the sampler chain has been properly thinned (see Theorem 2 below). We then split the samples into two subsets: a training set $\boldsymbol{\theta}^{1}=\{\theta_{1},\dots,\theta_{S}\}$ and a calibration set $\boldsymbol{\theta}^{2}=\{\theta_{S+1},\dots,\theta_{T}\}$. The key idea is to compute, for each calibration sample $\theta\in\boldsymbol{\theta}^{2}$, a pseudo-density score
$\theta\mapsto s(\theta;\boldsymbol{\theta}^{1})\in\mathbb{R},$
which is estimated using the training set $\boldsymbol{\theta}^{1}$ and is designed to capture a notion of likelihood or density of $\theta$ under the posterior distribution. This will be achieved by defining $s(\cdot;\boldsymbol{\theta}^{1})$ as a kernel density estimator (KDE) based on a discrepancy $\mathcal{D}:\Theta\times\Theta\to\mathbb{R}_{+}$ between parameters.
Once the scoring rule is available, the CBI framework proceeds through the following three main steps:
@@
    $\theta_{\star}\in\arg\max_{\theta\in\boldsymbol{\theta}^{2}}s(\theta;\boldsymbol{\theta}^{1}).$
    Because $s(\cdot;\boldsymbol{\theta}^{1})$ is a posterior pseudo-density function, $\theta_{\star}$ may be interpreted as a pseudo-maximum-a-posteriori (pseudo-MAP) estimator.
  • Step 2: credible region. By exchangeability of the calibration scores, standard conformal prediction arguments yield a credible region $\mathcal{C}_{1-\alpha}(\boldsymbol{\theta}^{2})\subseteq\Theta$, based on the quantiles of $\{s(\theta;\boldsymbol{\theta}^{1}):\theta\in\boldsymbol{\theta}^{2}\}$, having a prescribed $(1-\alpha)\times 100\%$ coverage under $\Pi$ (for $\alpha\in(0,1)$, in a precise statistical sense to be defined later). The key conceptual step leading to this construction is the reinterpretation of inference based on posterior Monte Carlo samples as prediction on the parameter space according to the posterior distribution, and to leverage conformal prediction tools accordingly. Two appealing features of this procedure are that (i) the attached coverage guarantee is independent of any feature of the target posterior $\Pi$, accommodating parameter spaces of arbitrary dimension and with almost no structure, and (ii) for a new sample $\theta$, it provides a simple way to define a standardized measure of typicality $\hat{p}(\theta;\boldsymbol{\theta}^{2})\in[0,1]$ under $\Pi$, which may also be formally interpreted as a $p$-value under an intuitive distributional null hypothesis.
  • Step 3: multimodality analysis. Finally, the interpretation of $\theta\mapsto s(\theta;\boldsymbol{\theta}^{1})$ as a posterior pseudo-density function allows one to explore the possibly multimodal structure of the posterior distribution through density-based clustering techniques. By equipping $\Theta$ with the discrepancy $\mathcal{D}$, one can identify distinct high-density regions, thereby revealing whether the posterior distribution places substantial mass around multiple, well-separated parameter values.
@@
While CBI can be described in full generality without reference to any particular statistical model, it is useful to ground the discussion in a concrete setting involving a parameter space $\Theta$ of interest. Therefore, before presenting a general treatment of CBI, in Section 2 we first introduce a fundamental class of models, namely random partition models, on which the strengths of our methodology can be clearly demonstrated. One advantage of working with random partition models is the existence of a rich literature addressing point estimation and uncertainty quantification based on posterior Monte Carlo samples over the space of data partitions (wade2018clustering, wade2023bayesian, balocchi2025understanding), providing a natural conceptual benchmark for our discussion. Moreover, since clustering is an inherently difficult problem (hennig2015true), Bayesian posteriors over data partitions often display substantial uncertainty, frequently in the form of multimodality, which our framework is specifically designed to handle. After this, in Sections 3, 4, and 5 we describe CBI in detail and illustrate it with simulated data analyzed with random partition models. [^2] However, see the Supplementary Material for two additional experiments applying CBI to models where the parameter space consists of discrete mixing measures endowed with the Wasserstein-1 distance (nguyen2013convergence, villani2008optimal) and covariance matrices endowed with the operator norm distance. Section 6 showcases two real-world applications of CBI to Bayesian clustering of (i) the classical Galaxy velocities dataset (roeder1990density), and (ii) event-related potential (ERP) waveform functional data (song2025repulsive, kappenman2021erp). Section 7 concludes the article with a discussion of the current limitations and future directions related to CBI.
@@ 2 Motivation: random partition models
allows for coherent inference on the clustering structure of the observed data. In practice, inference on clustering may be formulated at the latent level of a hierarchical density mixture model, where heterogeneity among subpopulations is interpreted as observations being generated from different components of a mixture model (fruhwirth2019handbook, lo1984class, escobar1995bayesian, lijoi2005hierarchical, barrios2013modeling, lijoi2020pitman), or, more recently, through approaches based on tailored losses (rigon2023generalized) or level sets (buch2024bayesian). We refer the reader to wade2023bayesian, grazian2023review, and the review section of balocchi2025understanding for comprehensive and up-to-date overviews of the field.
@@
The combinatorially large size of $\Theta$ typically prevents exact posterior evaluation, and approximate computational methods, most often in the form of MCMC sampling, are adopted instead. For instance, if the random partition model is specified within a hierarchical mixture framework, standard samplers (escobar1995bayesian, escobar1998computing, maceachern1998estimating, favaro2013mcmc, neal2000markov) will, at each iteration $t=1,\ldots,T$, draw a sample $\theta_{t}\sim\Pi$ (in addition to sampling the other mixture parameters). These samples are usually stored as vectors of cluster assignments $C_{t}=(c_{t1},\ldots,c_{tn})$, encoding the partition (equivalence) relation $\sim_{\theta_{t}}$ via $X_{i}\sim_{\theta_{t}}X_{j}$ if and only if $c_{ti}=c_{tj}$, for all $i,j=1,\ldots,n$. Assuming appropriate thinning of the chain and a warm start $\theta_{1}\sim\Pi$, the resulting samples $\theta_{1},\ldots,\theta_{T}$ will serve as the basis of our CBI procedure.
@@
A widely used metric on the space of partitions $\Theta$ is the Variation-of-Information (VI) metric (meila2007comparing), denoted by $\mathcal{D}_{VI}$ and defined as follows. Take any two partitions $\theta,\theta^{\prime}\in\Theta$ with associated cluster assignment vectors $C,C^{\prime}$ and respective numbers of clusters $K$ and $K^{\prime}$. Let $\{\bar{c}_{1},\ldots,\bar{c}_{K}\}$ and $\{\bar{c}^{\prime}_{1},\ldots,\bar{c}^{\prime}_{K^{\prime}}\}$ denote the sets of unique cluster labels in each vector, and define $n_{j,*}:=|\{i=1,\ldots,n:c_{i}=\bar{c}_{j}\}|$, $n_{*,k}:=|\{i=1,\ldots,n:c^{\prime}_{i}=\bar{c}^{\prime}_{k}\}|$, and $n_{j,k}:=|\{i=1,\ldots,n:c_{i}=\bar{c}_{j},\,c^{\prime}_{i}=\bar{c}^{\prime}_{k}\}|$, for all $j=1,\ldots,K$ and $k=1,\ldots,K^{\prime}$. Then,
$\mathcal{D}_{VI}(\theta,\theta^{\prime})$ $:=$ (1)
$-\,\sum_{j=1}^{K}$ $\frac{n_{j,*}}{n}\log_{2}\left(\frac{n_{j,*}}{n}\right)\,-\,\sum_{k=1}^{K^{\prime}}\frac{n_{*,k}}{n}\log_{2}\left(\frac{n_{*,k}}{n}\right)\,-\,2\sum_{j=1}^{K}\sum_{k=1}^{K^{\prime}}\frac{n_{j,k}}{n}\log_{2}\left(\frac{n\cdot n_{j,k}}{n_{j,*}\cdot n_{*,k}}\right).$
@@
meila2007comparing showed that $\mathcal{D}_{VI}$ is a proper metric on $\Theta$, and that it enjoys several appealing theoretical properties, such as its information-theoretic interpretation in terms of self-entropy and cross-entropy between partitions, while also being efficiently computable. Although alternative discrepancy measures have been proposed, such as the Binder loss (binder1978bayesian, dahl2006model, lau2007clustering), the adjusted Rand index (fritsch2009clustering), the normalized VI (vinh2010information, rastelli2018optimal), and more (quintana2003clustering, dahl2022search, nguyen2024summarizing), we focus on the VI metric in what follows, given its widespread use and well-established role as the standard metric in Bayesian clustering.
@@
The large size of $\Theta$ and its discrete nature make it an ideal example of a setting where point estimation and posterior uncertainty quantification are highly non-trivial, even when Monte Carlo samples from $\Pi$ are available. The popularity of clustering and the success of Bayesian approaches in this domain have given rise to an extensive literature on the topic. A seminal contribution is due to wade2018clustering, who popularized the use of the VI metric in Bayesian clustering and proposed to address point estimation via posterior expected VI minimization, while constructing (approximations of) VI credible balls with prescribed posterior mass for uncertainty quantification. [^3] See also the contemporaneous and independent preprint by page2025uncertainty. Computational and theoretical extensions of this framework, often but not exclusively focused on point estimation, include rastelli2018optimal, dahl2022search, nguyen2024summarizing, buch2024bayesian, among others. An insightful perspective was recently introduced by balocchi2025understanding, who emphasized the importance of accounting for multimodality in the posterior distribution over $\Theta$ and proposed summarizing $\Pi$ with the closest $L$-atom mixture of point masses in the Wasserstein-over-VI metric, yielding a fixed number $L$ of representative partitions corresponding to the distinct posterior modes.
@@
3 Conformalized Bayesian Inference: preliminaries and point estimation
Recall the problem we aim to solve: given iid posterior Monte Carlo samples $\boldsymbol{\theta}$, split into a training subset $\boldsymbol{\theta}^{1}$ and a calibration subset $\boldsymbol{\theta}^{2}$, how can we perform point estimation and uncertainty quantification in practice? Following the pipeline described in the introduction, the first key step is to define a scoring function $\theta\mapsto s(\theta;\boldsymbol{\theta}^{1})$ that captures a notion of pseudo-density (ferraty2006nonparametric) under the posterior distribution. We now propose a simple and robust solution.